Write a Pythagorean triplet whose one member is.   6
Write a Pythagorean triplet whose one member is.   6

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

For any natural number m, we know that 2m, m2–1, m2+1 is a Pythagorean triplet.
2m=6
⇒m=6/2=3
m^2–1=3^2–1=9–1=8
m^2+1= 3^2+1 = 9+1 = 10
∴(6, 8, 10) is a Pythagorean triplet.

​Thanks